# Neuromorphic detection and cooling of microparticles in arrays

### v3

@@
Micro-objects levitated in a vacuum are an exciting platform for precision sensing due to their low dissipation motion and the potential for control at the quantum level. Arrays of such sensors would offer increased sensitivity, directionality, and in the quantum regime the potential to exploit correlation and entanglement. We use neuromorphic detection via a single event based camera to record the motion of an array of levitated microspheres. We present a scalable method for arbitrary multiparticle tracking and control by implementing real-time feedback to cool the motion of three objects simultaneously, the first demonstration of neuromorphic sensing for real-time control at the microscale.
@@ Introduction
Modern technology relies on mechanical sensors, from accelerometers in mobile devices [Shaeffer2013] to wearable health monitors [Chircov2022]. As sensors are miniaturized, their surface-to-volume ratio increases and they dissipate more energy via their thermal contacts and through surface strain [Imboden2014], limiting their performance. By levitating nano- or micro-particles under ultra-high vacuum conditions, using optical, electrical or magnetic fields [Millen2020rev, G-Ballestero2021rev], one creates a mechanical oscillator with remarkably low dissipation [dania2024ultrahigh, Moore2021rev]. Force sensitivities of yoctonewtons [Liang2022, ranjit2016zeptonewton] and torque sensitivities at the level [ahn2020ultrasensitive] have been achieved, with levitating sensors achieving quality factors in excess of [dania2024ultrahigh], motivating researchers to use these systems to search for dark matter [afek2022coherent, yin2022experiments, Moore2021rev] and gravitational waves [arvanitaki2013detecting, Winstone2022].
@@
The control of levitated particles allows the exploration of a wide range of fundamental science [Millen2020rev, G-Ballestero2021rev], and the demonstration of cooling to the ground state of an optical potential [delic2020cooling, tebbenjohanns2021quantum, magrini2021real, kamba2022optical, Piotrowski2023, ranfagni2022two] opens the door to macroscopic quantum physics [millen2020quantum, barker2010, chang2010, romeroisart2010]. An emerging frontier in this field is the study of arrays of particles, which in the quantum regime would allow generation of entanglement [rudolph2022force] and tests of quantum gravity [bose2017spin, marletto2017gravitationally]. Interactions have been observed between pairs of levitated nanoparticles in optical [arita2018optical, rieser2022tunable, livska2023cold, penny2023sympathetic], electrodynamic (Paul) [bykov20233d] and magnetic [Slezak2019] traps. Detecting and controlling multiple particles in vacuum has so far involved either single particle control with sympathetic cooling [arita2022all, penny2023sympathetic, bykov20233d] or small arrays of optical traps [vijayan2023scalable, Hupfl2023]. Some applications will require the control of arrays of tens, or even thousands, of levitated particles [afek2022coherent].
@@
We use neuromorphic imaging for the control of arbitrary particle arrays across a wide field-of-view. Neuromorphic sensors are highly efficient detectors which mimic neurobiological information gathering [Ambrogio2018, roy2019towards, Xiao2020, Liao2021, Sharma2024]. Dynamic vision sensors (DVS) are neuromorphic sensors which mimic the retinal response [Vanarse2016], detecting changes across a threshold on each pixel in an array asynchronously to produce a stream of events [lichtsteiner2008128] ideally suited for object tracking [Ni2015]. Together with event-driven processing algorithms [nguyen2019real, rebecq2019high, cannici2019asynchronous], DVS can achieve microsecond temporal resolution, sub-millisecond latency and high dynamic range detection (>) with minimal data output at low power consumption [roy2019towards, gallego2020event, Sharma2024]. Therefore, DVS are highly suited to high-speed and real-time applications requiring low-power in environments with uncontrolled light levels such as in robotics [zhu2023neuromorphic], autonomous driving [gehrig2024low] and space flight [Sikorski2021], as well as finding applications in microscopy [Ni2012, mangalwedhekar2023achieving] and astronomy [Ng2022]. In this work, we use an event based camera (EBC) with integrated DVS to monitor the motion of an array of levitated particles with a bandwidth high enough to demonstrate real-time simultaneous feedback control of multiple particles in an array.
@@
We implement cold damping feedback [Poggio2007] to cool the motion of the levitated particles [Tebbenjohanns2019, Conangla2019], a technique with demonstrated quantum ground state cooling capabilities [tebbenjohanns2021quantum, magrini2021real, kamba2022optical]. We cool a single microsphere to sub-Kelvin temperatures and single degrees-of-freedom of multiple particles. In this work, the number of objects we can simultaneously control is only limited by the number of output channels in our feedback electronics. This single-device method for cooling and controlling particles in arrays is readily scalable due to the low data output of neuromorphic detection. Arrays of cooled micro-sensors will lead to enhanced signal-to-noise sensing through sensor fusion [Palit1997, Sasiadek2002, Nazarahari2021], enable force gradient sensing [rudolph2022force], and provide a larger interaction area without increasing the mass of the sensor [afek2022coherent]. Due to the low power consumption of neuromoprhic detectors our presented methods are ideal for integration into chip-scale technology [melo2024vacuum].
@@ Results
We levitate arrays of charged diameter silica microspheres in a linear Paul trap under vacuum conditions [Kane2010, Kuhlicke2014, Millen2015, Delord2017] (see Methods) and record their motion using an event based camera (EBC) [ren2022event], see Fig. 1(a). The charged particles form a stable array due to the Coulomb repulsion between them. Our particular Paul trap geometry (Fig. 1(b)), and the particles’ distribution of charge, means that our particle arrays are non-uniform.
The EBC uses a neuromorphic DVS, which is an array of independent pixels featuring contrast detectors which output an event [roy2019towards] in response to light levels on the pixel crossing a user-defined threshold. Pixels which do not experience the required level-change output no signal, removing the data-redundancy present in conventional cameras [roy2019towards], while allowing the full sensor to be used at all times, sometimes referred to as a dynamic region-of-interest. The EBC hardware bundles asynchronous events into equal-length frames, and uses filters to identify objects within its field-of-view [Ni2015], afterwhich a proprietary generic tracking algorithm (GTA) tracks the motion of each object independently. We have previously demonstrated single-particle object tracking with an SNR above 35 dB, and for a more detailed analysis of EBC performance in the context of levitated microparticles see [ren2022event].
The EBC allows the tracking of multiple objects with a high bandwidth and a linear scaling in data output with the number of tracked objects, as compared to a rapid increase in data volume with increased region-of-interest in a conventional camera. In our system, with fixed magnification, tracking a single particle at 1 kHz using the entire field-of-view (mm2) of the EBC uses kB s-1 as compared to 64,800 kB s-1 using a standard CMOS camera (Thorlabs CS165MU/M) at the same frame rate and field-of-view. This means that the EBC can track many hundreds of particles before the data volume becomes comparable to standard camera technology. By not having to restrict the region-of-interest, the EBC can track objects dispersed over several hundred micrometres whilst retaining high spatial resolution [ren2022event]. The low data volume leads to a correspondingly low power consumption, see Supplemental Materials S5 for more details.
@@
In Fig. 2(a) we reconstruct the motion and relative position of four levitated particles obtained from the 1 kHz tracking algorithm of the EBC, which can be accessed in real-time. Particles levitated by a Paul trap undergo harmonic motion. In Fig. 2(b) we generate the power spectral density (PSD) from the motion of each particle in the array of four to analyze their motion in frequency space. Each particle has a different charge-to-mass ratio and is levitated in a different part of the confining field, meaning each particle has different frequency modes of oscillation, which are well separated under vacuum conditions. Below, we use this fact to independently cool the motion of multiple particles. Coulomb interactions between the particles are evident [Slezak2019, Svak2021, penny2023sympathetic, bykov20233d], as shown by the collective modes in Fig. 2(b), see Supplementary Materials S3. The particles are not coupled in the direction, which is the mode we use for multiparticle cooling. The signal-to-noise for the different particles varies due to non-uniform illumination and varying coupling to electronic noise. All data in this work is taken at gas pressures of 2.0-mbar unless otherwise stated.
@@
There are many reasons why it is desirable to control the energy of a levitated particle. Although the sensitivity of a levitated sensor does not increase through cooling [monteiro2020force], rapid damping of the motion increases the stability and measurement bandwidth of the system. Reduction of the particle energy to the ground-state of the levitating potential [delic2020cooling, magrini2021real, piotrowski2023simultaneous] opens up a toolbox of quantum control [G-Ballestero2021rev] and sensitivity enhancement [Mason2019, Cernotik2020] mechanisms.
Cold damping is a feedback method whereby a force proportional to the velocity of an oscillator opposes its motion. Depending on the phase of the feedback force relative to the motion, this method can cause damping or amplification of the particle motion without adding additional noise, hence the terminology “cold damping”. In reality, input and output noise of the feedback electronics still limits cooling.
@@
Figure 3(a) shows the PSD of a particle’s motion along the axis as it is cooled via cold damping. The shape of the position PSD is given by [Conangla2019]:
@@
where , is the particle’s mass, is the mode frequency, is the bath temperature and is the total momentum damping rate on the particle’s motion, with being the calculable momentum damping rate due to the pressure of the surrounding gas [Millen2020rev] and being the additional damping controlled by the feedback gain, which is feedback-phase dependent. The terms and are the feedback circuit noise and detector noise respectively, and are modeled as having a constant spectral density. When is large the feedback can introduce extra noise and leads to heating, as discussed in [Conangla2019, penny2021performance, melo2024vacuum]. The parameters , and can be extracted from a measured PSD by fitting equation (Single particle cold damping using neuromorphic imaging) to the data. Due to voltage noise from the amplifiers driving our Paul trap, the equilibrium temperature of our particles without cooling ranges from , depending on their charge and spatial location in the trap, hence we express temperatures as a ratio.
Experimentally, the temperature of each mode can be extracted from the integral of the measured PSD over the corresponding resonance peak [Conangla2019]. According to the equipartition theorem, the temperature of a levitated particle experiencing cold damping is [Conangla2019, penny2023sympathetic]:
@@
In Fig. 3(b) we show the effect of increasing on the temperature of a single mode of a single particle, and compare the temperature measured using the integral of the PSD to the model in equation (2), with as a free parameter and the other parameters extracted from the PSDs in Fig. 3(a) using equation (Single particle cold damping using neuromorphic imaging).
The cooling depends on the phase between the feedback signal, which is proportional to velocity, and the detected motion of the particle, with the temperature varying as [dania2021optical]:
@@
where is the uncontrollable phase delay caused by electronics and data processing, and is a controllable phase delay generated by our feedback electronics. Equation (3) is fit to experimental data in Fig. 3(c), with and as free parameters. The fitted value of is , noting that one period of phase delay does not significantly effect cooling for an underdamped oscillator [debiossac2020], see Supplementary Materials S4. The value of obtained by fitting the data in Fig. 3(c) with equation (3) agrees with the value obtained at the same feedback gain by fitting the data in Fig. 3(a) with equation (Single particle cold damping using neuromorphic imaging), (Hz, Hz respectively).
@@
Finally, we combine the optimized and to push cooling to the limit by reducing through reducing the gas pressure in the experiment, as shown in Fig. 3(d). At low pressures we reach the noise floor of our system, indicated by the grey region, at a temperature corresponding to , representing -30 dB of cooling. To further improve cooling one can improve particle illumination and imaging, decrease the noise in the levitation electronics, and replace the GTA of the EBC with an optimized tracking algorithm. Object tracking has the potential to track levitated microparticles at the shot-noise limit [Werneck2024].
@@
We optimize and fix the feedback parameters, then lower the background pressure to reduce , hence lowering the temperature . The temperature is limited not by our noise floor, but by imperfect filtering pumping energy from the feedback signal into the mode, since it is close in frequency to the mode. At low pressures this causes the particle to become unstable, preventing us from further lowering the pressure.
@@
In Fig. 4(c) we cool the mode of three different particles, with the corresponding PSDs shown in Fig. 4(d). To the best of our knowledge, this is the first demonstration of cooling more than two levitated particles. The issue of imperfect filtering is more pronounced when dealing with more particles, as there are more modes of the system overlapped with the unfiltered noise. We still achieve better than 7 dB of cooling. Filtering can be improved either by separating the particle modes and applying band-pass filtering, or through the use of phase locked loops [Jain2016, Vovrosh2017]. For the data in Figs. 4(b-d) we adjust the particle spacing via the Paul trap voltages until the Coulomb interaction is weak enough such that there are no coupled modes, see Supplementary Materials S2. Collective modes in particle arrays can be cooled via sympathetic cooling [penny2023sympathetic, bykov20233d].
The EBC used in this study has a sensor size of pixels, with each particle image occupying pixels (the coloured boxes in Fig. 1(c)) and having a motional amplitude of 4 pixels. As long as the centre of the particles are separated by approximately 60 pixels, the particles can be individually tracked. Hence, without changing our imaging system we could simultaneously track of order 500 particles with this EBC. Considering the fact that object tracking allows for sub-pixel resolution [Bullier2019, Ni2012], by changing the magnification of the imaging system this EBC would be capable of simultaneously tracking at least 2000 levitated microparticles, with a correspondingly high data volume.
@@ Discussion
We have presented a scalable method for the detection and control of microparticles levitated in an array using a single neuromorphic detector. Neuromorphic imaging is ideally suited to this task, due to its natural affinity with detecting the motion of multiple objects [Ni2015, Apps2025] and low data-transfer rate [roy2019towards]. The tracking speed in our work is limited by the proprietary tracking algorithm of the EBC. Commercial neuromorphic imaging sensors, such as the one used in this study, transfer data from the sensor to the camera hardware at GHz rates [gallego2020event]. The development of custom algorithms has enabled object tracking at 30 kHz by working with the asychronous data streamed from a DVS using only 4MB of RAM on a standard 2.9 GHz Dual Core CPU [Ni2012]. For a fixed frame rate the data volume is fixed regardless of the particle motion frequency. If the frame-rate is increased data is transferred more rapidly from the camera and the data volume and bandwidth increases. By pushing above 100 kHz, neuromorphic sensors would be suitable for feedback cooling optically levitated particles to the quantum ground state of motion [magrini2021real], considering the shot-noise limited potential of object tracking [Bullier2019, Werneck2024]. This would require custom tracking algorithms and interfacing the sensor directly with FPGA or neuromorphic processing electronics [roy2019towards, Furmonas2022], which would also enable the read-out and control of object alignment and rotation [Kim2016, Apps2025].
@@
We believe that the particle control method presented in this work could be extended to an array of order 100 microparticles. Multichannel FPGA systems with high-quality digital filters are a common tool in research labs. Paul traps are stable at low pressures [bykov2019direct], where the motional frequencies of levitated particles have sub-Hz linewidths [pontin2020ultranarrow]. The naturally varying charge-to-mass ratio of charged microparticles, along with application of electric field gradients, will enable the spectral separation of motional modes, making possible single-particle control and cooling even for large arrays. The neuromorphic detection and cooling presented here is independent of the levitation method, as long as there is optical illumination, meaning it is suitable for small dielectrics in optical traps, charged absorptive materials in Paul traps (such as organic material) [Conangla2019], and magnetically levitated objects.
@@
Since the motion of levitated sensors is well understood, simple machine-learning could be used to optimize all of the feedback parameters [Conangla2019] in an array, and optimal tracking algorithms used which have been shown to enable quantum-level control [magrini2021real]. When combined with the low power-consumption of neuromorphic imaging technology (less than 30 mW per tracked particle, see Supplementary Materials S5), and great progress in chip-scale particle levitation [melo2024vacuum], integrated devices containing arrays of quantum sensors are closer to being a reality.

### v4

@@
Micro-objects levitated in a vacuum are an exciting platform for precision sensing due to their low dissipation motion and the potential for control at the quantum level. Arrays of such sensors would offer increased sensitivity, directionality, and in the quantum regime the potential to exploit correlation and entanglement. We use neuromorphic detection via a single event based camera to record the motion of an array of levitated microspheres. We present a scalable method for arbitrary multiparticle tracking and control by implementing real-time feedback to simultaneously cool the motion of three uncoupled microscale objects.
@@ Introduction
Modern technology relies on mechanical sensors, from accelerometers in mobile devices [1] to wearable health monitors [2]. As sensors are miniaturized, their surface-to-volume ratio increases and they dissipate more energy via their thermal contacts and through surface strain [3], limiting their performance. By levitating nano- or micro-particles under ultra-high vacuum conditions, using optical, electrical or magnetic fields [4, 5], one creates a mechanical oscillator with remarkably low dissipation [6, 7]. Force sensitivities of yoctonewtons [8, 9] and torque sensitivities at the level [10] have been achieved, with levitating sensors achieving quality factors in excess of [6], motivating researchers to use these systems to search for dark matter [11, 12, 7] and gravitational waves [13, 14].
@@
The control of levitated particles allows the exploration of a wide range of fundamental science [4, 5], and the demonstration of cooling to the ground state of an optical potential [15, 16, 17, 18, 19, 20] opens the door to macroscopic quantum physics [21, 22, 23, 24]. An emerging frontier in this field is the study of arrays of particles, which in the quantum regime would allow generation of entanglement [25] and tests of quantum gravity [26, 27]. Interactions have been observed between pairs of levitated nanoparticles in optical [28, 91, 30, 87], electrodynamic (Paul) [32] and magnetic [33] traps. Detecting and controlling multiple particles in vacuum has so far involved either single particle control with sympathetic cooling [34, 87, 32] or small arrays of optical traps [35, 36]. Some applications will require the control of arrays of tens, or even thousands, of levitated particles [11].
@@
We use neuromorphic imaging for the control of arbitrary particle arrays across a wide field-of-view. Neuromorphic sensors are highly efficient detectors which mimic neurobiological information gathering [37, 38, 39, 40, 41]. Dynamic vision sensors (DVS) are neuromorphic sensors which mimic the retinal response [42], detecting changes across a threshold on each pixel in an array asynchronously to produce a stream of events [43] ideally suited for object tracking [44]. Together with event-driven processing algorithms [45, 46, 47], DVS can achieve microsecond temporal resolution, sub-millisecond latency and high dynamic range detection (>) with minimal data output at low power consumption [38, 48, 41]. Therefore, DVS are highly suited to high-speed and real-time applications requiring low-power in environments with uncontrolled light levels such as in robotics [49], autonomous driving [50] and space flight [51], as well as finding applications in microscopy [52, 53] and astronomy [54]. In this work, we use an event based camera (EBC) with integrated DVS to monitor the motion of an array of levitated particles with a bandwidth high enough to demonstrate real-time simultaneous feedback control of multiple uncoupled particles in an array.
@@
We implement cold damping feedback [94] to cool the motion of the levitated particles [56, 95], a technique with demonstrated quantum ground state cooling capabilities [16, 17, 18]. We cool a single microsphere to a temperature of a few Kelvin and single degrees-of-freedom of multiple particles. In this work, the number of objects we can simultaneously control is only limited by the number of output channels in our feedback electronics. This single-device method for cooling and controlling particles in arrays is readily scalable due to the low data output of neuromorphic detection. Arrays of cooled micro-sensors will lead to enhanced signal-to-noise sensing through sensor fusion [58, 59, 60], enable force gradient sensing [25], and provide a larger interaction area without increasing the mass of the sensor [11]. Due to the low power consumption of neuromorphic detectors our presented methods are ideal for integration into chip-scale technology [93].
@@ Results
We levitate arrays of charged m diameter silica microspheres in a linear Paul trap under vacuum conditions [62, 63, 64, 65] (see Methods) and record their motion using an event based camera (EBC) [89], see Fig. 1(a). The charged particles form a stable array due to the Coulomb repulsion between them. Our particular Paul trap geometry (Fig. 1(b)), and the particles’ distribution of charge, means that our particle arrays are non-uniform.
The EBC uses a neuromorphic DVS, which is an array of independent pixels featuring contrast detectors which output an event [38] in response to light levels on the pixel crossing a user-defined threshold. Pixels which do not experience the required level-change output no signal, removing the data-redundancy present in conventional cameras [38], while allowing the full sensor to be used at all times, sometimes referred to as a dynamic region-of-interest. The EBC hardware bundles asynchronous events into equal-length frames, and uses filters to identify objects within its field-of-view [44], after which a proprietary generic tracking algorithm (GTA) tracks the motion of each object independently. We have previously demonstrated single-particle object tracking with an SNR above 35 dB, and for a more detailed analysis of EBC performance in the context of levitated microparticles see [89].
The EBC allows the tracking of multiple objects with a high bandwidth and a linear scaling in data output with the number of tracked objects, as compared to a rapid increase in data volume with increased region-of-interest in a conventional camera. In our system, with fixed magnification, tracking a single particle at 1 kHz using the entire field-of-view (mm2) of the EBC uses kB s-1 as compared to 64,800 kB s-1 using a standard CMOS camera (Thorlabs CS165MU/M) at the same frame rate and field-of-view. This means that the EBC can track many hundreds of particles before the data volume becomes comparable to standard camera technology. By not having to restrict the region-of-interest, the EBC can track objects dispersed over several hundred micrometres whilst retaining high spatial resolution [89]. The low data volume leads to a correspondingly low power consumption, see Supplemental Materials S5 for more details.
@@
In Fig. 2(a) we reconstruct the motion and relative position of four levitated particles obtained from the 1 kHz tracking algorithm of the EBC, which can be accessed in real-time. Particles levitated by a Paul trap undergo harmonic motion. In Fig. 2(b) we generate the power spectral density (PSD) from the motion of each particle in the array of four to analyze their motion in frequency space. Each particle has a different charge-to-mass ratio (see Supplementary Materials S3) and is levitated in a different part of the confining field, meaning each particle has different frequency modes of oscillation, which are well separated under vacuum conditions. Below, we use this fact to independently cool the motion of multiple particles.
The four particles are aligned along the axis. Motion of the charged particles in this direction leads to coupling between them via the Coulomb interaction [33, 67, 87, 32]. This is evident via the collective modes seen in Fig. 2(b). By controlling the separation between the particles, we can control the coupling strength, see Supplementary Materials S2 for further details. For the multi-particle cooling presented below, we work in a regime where the coupling between the modes is too small to measure, and perform cooling along the direction.
The signal-to-noise for the different particles varies due to non-uniform illumination and varying coupling to electronic noise. It ranges from 3-30, with exact values given in Supplementary Materials S3. All data in this work is taken at gas pressures of 2.0-mbar unless otherwise stated.
@@
There are many reasons why it is desirable to control the energy of a levitated particle. Although the sensitivity of a levitated sensor does not increase through cooling [68], rapid damping of the motion increases the stability and measurement bandwidth of the system. Reduction of the particle energy to the ground-state of the levitating potential [15, 17, 69] opens up a toolbox of quantum control [5] and sensitivity enhancement [70, 71] mechanisms.
Cold damping is a feedback method whereby a force proportional to the velocity of an oscillator opposes its motion. Depending on the phase of the feedback force relative to the motion, this method damps (cools) or amplifies (heats) the oscillator. When cooling, input and output noise of the feedback electronics limits the ultimate temperature.
@@
Figure 3(a) shows the PSD of a particle’s measured motion along the axis as it is cooled via cold damping. The shape of the PSD from the in-loop detector is given by [94, 95]:
@@
where , is the particle’s mass, is the mode frequency, is the bath temperature and is the total momentum damping rate on the particle’s motion, with being the calculable momentum damping rate due to the pressure of the surrounding gas [4] and being the additional damping controlled by the feedback gain, which is feedback-phase dependent, with the uncontrollable phase delay caused by electronics and data processing. The term is the feedback circuit noise that is modeled as white noise. When is large the feedback can introduce extra noise and leads to heating, as discussed in [95, 72, 93]. The parameters , and can be extracted from a measured PSD by fitting equation (Single particle cold damping using neuromorphic imaging) to the data. Due to voltage noise from the amplifiers driving our Paul trap, the equilibrium temperature of our particles without cooling ranges from , depending on their charge and spatial location in the trap, hence we express temperatures as a ratio. The equilibrium temperatures for all particles in this paper are given in Supplementary Material S6.
According to the equipartition theorem, the temperature of a levitated particle experiencing cold damping is [95, 87, 73]:
@@
Experimentally, the temperature of each mode can be extracted from equation (2) after fitting the measured PSD over the corresponding resonance peak with equation (Single particle cold damping using neuromorphic imaging) [95]. In Fig. 3(b) we show the effect of increasing on the temperature of a single mode of a single particle.
@@
The cooling depends on the phase between the feedback signal, which is proportional to velocity, and the detected motion of the particle. Equation (2) is fit to experimental data as is varied in Fig. 3(c), with and as free parameters. The fitted value of is , noting that one period of phase delay does not significantly effect cooling for an underdamped oscillator [92], see Supplementary Materials S4. The value of obtained by fitting the data in Fig. 3(c) with equation (2) agrees with the value obtained at the same feedback gain by fitting the data with equation (Single particle cold damping using neuromorphic imaging), (Hz, Hz respectively.
@@
Finally, in Fig. 3(d) we show the variation in temperature with by reducing the gas pressure, with and fixed at around Hz and , respectively. At a pressure of mbar, we reach the noise floor of our system, indicated by the grey region, at a temperature corresponding to , representing 17 dB of cooling. The optimal can be derived from equation (2), see Supplementary Materials S7. To further improve cooling one can improve particle illumination and imaging, decrease the noise in the levitation electronics, and replace the GTA of the EBC with an optimized tracking algorithm. Object tracking has the potential to track levitated microparticles at the shot-noise limit [75].
@@
We optimize and fix the feedback parameters, then lower the background pressure to reduce , hence lowering the temperature . The temperature is limited not by our noise floor, but by imperfect filtering pumping energy from the feedback signal into the mode, since it is close in frequency to the mode. At low pressures this causes the particle to become unstable, preventing us from further lowering the pressure. Cross-talk between particle modes would also limit the ultimate cooling temperature, which could be resolved by better design of the feedback electrodes [76].
@@
In Fig. 4(c) we cool the mode of three different particles, with the corresponding PSDs shown in Fig. 4(d). The issue of imperfect filtering is more pronounced when dealing with more particles, as there are more modes of the system overlapped with the unfiltered noise. We still achieve better than 7 dB of cooling. Filtering can be improved either by separating the particle modes and applying band-pass filtering, or through the use of phase locked loops [77, 78]. For the data in Figs. 4(b-d) we adjust the particle spacing via the Paul trap voltages until the Coulomb interaction is weak enough such that there are no coupled modes, see Supplementary Materials S2. Collective modes in particle arrays can be cooled via sympathetic cooling in the limit that the feedback damping rate is smaller than the coupling strength between the modes [87, 32].
The EBC used in this study has a sensor size of pixels, with each particle image occupying pixels (the coloured boxes in Fig. 1(c)) and having a motional amplitude of 4 pixels. As long as the centre of the particles are separated by approximately 60 pixels, the particles can be individually tracked. Hence, without changing our imaging system we could simultaneously track of order 500 particles with this EBC. Considering the fact that object tracking allows for sub-pixel resolution [79, 52], by changing the magnification of the imaging system this EBC would be capable of simultaneously tracking at least 2000 levitated microparticles, with a correspondingly high data volume.
@@ Discussion
We have presented a scalable method for the detection and control of microparticles levitated in an array using a single neuromorphic detector. Neuromorphic imaging is ideally suited to this task, due to its natural affinity with detecting the motion of multiple objects [44, 80] and low data-transfer rate [38]. The tracking speed in our work is limited by the proprietary tracking algorithm of the EBC. Commercial neuromorphic imaging sensors, such as the one used in this study, transfer data from the sensor to the camera hardware at GHz rates [48]. The development of custom algorithms has enabled object tracking at 30 kHz by working with the asynchronous data streamed from a DVS using only 4MB of RAM on a standard 2.9 GHz Dual Core CPU [52]. For a fixed frame rate the data volume is fixed regardless of the particle motion frequency. If the frame-rate is increased data is transferred more rapidly from the camera and the data volume and bandwidth increases. By pushing above 100 kHz, neuromorphic sensors would be suitable for feedback cooling optically levitated particles to the quantum ground state of motion [17], considering the shot-noise limited potential of object tracking [79, 75]. This would require custom tracking algorithms and interfacing the sensor directly with FPGA or neuromorphic processing electronics [38, 81], which would also enable the read-out and control of object alignment and rotation [82, 80].
@@
We believe that the particle control method presented in this work could be extended to an array of order 100 microparticles, (see Supplementary Materials S8). Multichannel FPGA systems with high-quality digital filters are a common tool in research labs. Paul traps are stable at low pressures [83], where the motional frequencies of levitated particles have sub-Hz linewidths [84]. The naturally varying charge-to-mass ratio of charged microparticles, along with application of electric field gradients, will enable the spectral separation of motional modes, making possible single-particle control and cooling even for large arrays. The neuromorphic detection and cooling presented here is independent of the levitation method, as long as there is optical illumination, meaning it is suitable for small dielectrics in optical traps, charged absorptive materials in Paul traps (such as organic material) [95], and magnetically levitated objects.
@@
Since the motion of levitated sensors is well understood, simple machine-learning could be used to optimize all of the feedback parameters [95] in an array, and optimal tracking algorithms used which have been shown to enable quantum-level control [17]. When combined with the low power-consumption of neuromorphic imaging technology (less than 30 mW per tracked particle, see Supplementary Materials S5), and great progress in chip-scale particle levitation [93], integrated devices containing arrays of quantum sensors are closer to being a reality.